OpenStudy (anonymous):

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer): 7 cos θ sin θ + 2 cos θ = 0

OpenStudy (anonymous):

I tried it and got \[\theta=\frac{ \Pi }{ 2 } and \frac{ 3\Pi }{ 2 }\] ...but its wrong

OpenStudy (slaaibak):

First factor out the cos theta: \[\cos(\theta)(7\sin(\theta) + 2) =0\] \[\rightarrow \cos \theta =0 \] or \[7\sin(\theta) + 2 = 0\] Jut solve those two.

OpenStudy (anonymous):

yeah, that's what I did and since the second one you cant find the sine of -2/7 I can't solve for that one. Which leaves only pi/2 and 3pi/2 for the first one.

OpenStudy (campbell_st):

you can find sin(x) = 2/7 then use the angles in any quadrant to fin the solution sin is negative in the 3rd (pi + x) and 4th quadrant (2pi - x) this will allow you to find all the solutions.